Identify the system of equations used to solve the following problem. A laboratory employee is mixing a 10% saline solution with 4% saline solution. How much of each solution is needed to make 500 milliliters of a 6% solution?
Answer by fractalier(6550)

The equation you need is (letting x equal the amount of 10% saline solution)
x(10%) + (500-x)(4%) = 500(6%) or
x(.1) + (500-x)(.04) = 500(.06)